A cylindrical container with a cross-sectional area of 61.2 cm2 holds a fluid of density 816 kg/m3. At the bottom of the container the pressure is 120 kPa. Assume Pat = 101 kPa. Depth of the fluid = h=2.37m


Find the pressure at the bottom of the container after an additional 2.05×10−3 m3 of this fluid is added to the container. Assume that no fluid spills out of the container.
Answer #1

Depth of fluid added h2 =V/A = 2.05×10-3 / (61.2×10-4) =0.335 m

Pressure at the bottom P = Po + Dg(h1+h2)

P = 101000+(816×9.81×(2.37+0.335))

P = 122.6 kPa